OpenStudy (anonymous):

Which expression is equivalent to 16^3?

OpenStudy (astrophysics):

What are the options

OpenStudy (anonymous):

2^11

OpenStudy (anonymous):

2^7

OpenStudy (anonymous):

2^12

OpenStudy (anonymous):

2^64

OpenStudy (astrophysics):

Keep simplifying, can you simplify 16 in exponents?

OpenStudy (astrophysics):

\[16=4^2\] right?

OpenStudy (anonymous):

no

OpenStudy (astrophysics):

why not?

OpenStudy (anonymous):

none of those are the option

OpenStudy (astrophysics):

Remember \[4^2 = 4 \times 4\] the exponent just tells you how many times you multiply by itself

OpenStudy (astrophysics):

So we have now \[(4^2)^3 = 4^6\] now can we simplify 4?

OpenStudy (anonymous):

so itll be 2^64

OpenStudy (astrophysics):

No, \[4 = 2^2 \]

OpenStudy (astrophysics):

\[(2^2)^6\]

OpenStudy (anonymous):

im so confused

OpenStudy (astrophysics):

Ok so we have \[16^3 \] and we need it to be one of the options right

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

So we just simplify the integer we have, as it goes into 2, we are trying to get that 2 first

OpenStudy (anonymous):

okay

OpenStudy (astrophysics):

\[16^3 = (4^2)^3 = 4^{2 \times 3} = 4^6 \] see how that works?

OpenStudy (astrophysics):

Because \[4^2 = 16\]

OpenStudy (astrophysics):

\[4^2 = 4 \times 4 = 16\]